OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

perpendicular tells you what about the slope?

OpenStudy (anonymous):

Uhh....I don't know...Im really bad at math. I dont know what slope isss

OpenStudy (anonymous):

slope is m in the equation of y=mx+b

OpenStudy (anonymous):

y = mx + b where: m is the slope b is the y-intercept.

OpenStudy (anonymous):

Idk..I need hellpp.

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm still confusedd.. Smhh aha

OpenStudy (anonymous):

Perpendicular means the negative reciprocal of the slope, for example 2, so it would be -1/2

OpenStudy (anonymous):

y = 2x + 2 and passes through the point (6, 3). This line has a slope = 2 Thus a perpendicular line will have a slope = -1/2

OpenStudy (anonymous):

So you can ELIMINATE choice A & D

OpenStudy (anonymous):

Is it b or c?

OpenStudy (anonymous):

Yes, of course if you ELIMINATE A and D, B and C are left, duh.

OpenStudy (anonymous):

I know. I said is it b or c....Like which one is it..

jimthompson5910 (jim_thompson5910):

the line must go through (6,3) so plug in x = 6 and you *should* get y = 3 if you don't get y = 3, then (6,3) does NOT lie on the line

OpenStudy (anonymous):

there you go @jim_thompson5910 got it :)

OpenStudy (anonymous):

the answer is potato XD nah is C y=-0.5*6 + 6=-3+6=3